OpenStudy (anonymous):

The equation y = 16x + 560 describes the number of calories in a recipe, where y is the total number of calories and x is the number of teaspoons of sugar added. The table shows the number of calories in another recipe, where y is the total number of calories and x is the number of teaspoons of honey added. How many calories are in a teaspoon of sugar, and how many are in a teaspoon of honey? Teaspoons of honey 1 2 4 7 Total calories 504 568 696 888 A teaspoon of sugar has calories. A teaspoon of honey has calories.

OpenStudy (anonymous):

@sleepyjess

OpenStudy (sleepyjess):

Well, it gives you the calories for honey, and for sugar, I think you just substitute in 1 for x in the equation.

OpenStudy (sleepyjess):

Well, maybe, now I'm not sure after reading it again

OpenStudy (anonymous):

so how do i put in a equation

OpenStudy (sleepyjess):

Maybe @Directrix could help?

Directrix (directrix):

Let me think @sleepyjess Sorry about that trig problem.

Directrix (directrix):

Did anybody get the equation for the data in the table?

OpenStudy (sleepyjess):

Let me work it out really quick

OpenStudy (anonymous):

it wants to know how many teaspoons of honey and sugar

OpenStudy (anonymous):

35 ?

Directrix (directrix):

@sleepyjess What did you get for the calories/honey equation?

OpenStudy (sleepyjess):

I pmed it to you

OpenStudy (anonymous):

i got wrong thoe all they way i got 0%

OpenStudy (anonymous):

on my whole quiz

Directrix (directrix):

Does anybody agree with this for the calories/honey: The equation of the line that passes through the points (1,504) and (2,568) is y=64x+440

OpenStudy (sleepyjess):

@hartnn

OpenStudy (sleepyjess):

I came up with the same equation @Directrix

Directrix (directrix):

y = 16x + 560 calories/sugar y=64x+440 calories/honey

hartnn (hartnn):

a teaspoon of honey= 1 teaspoon of honey = 504 cal

hartnn (hartnn):

a teaspoon of sugar = 1 teaspoon of sugar = 560+16 = 576 cal

OpenStudy (sleepyjess):

I wasn't sure about the calories of sugar though because it says "y is the total number of calories and x is the number of teaspoons of sugar added." not y is the calories of sugar

hartnn (hartnn):

x teaspoon = y calories

OpenStudy (sleepyjess):

That is what I was originally thinking, but then I started to question myself :/

hartnn (hartnn):

y = 16x + 560 calories y=64x+440 calories are correct

Directrix (directrix):

The intersection is (5/2, 600) which I am now confused over the meaning of it.

Directrix (directrix):

x is the number of teaspoons, y the calories. 5/2 teaspoons of either sweetener produce the same calories. But, that is not the question.

hartnn (hartnn):

why is intersection requireD? a teaspoon of honey= 1 teaspoon of honey = 504 cal a teaspoon of sugar = 1 teaspoon of sugar = 560+16 = 576 cal thats it!

Directrix (directrix):

Alrighty, then. I need to cogitate on this. Thanks, @hartnn

Directrix (directrix):

And, thanks @sleepyjess and @Lolafire

OpenStudy (girlgoyleh.):

the answer was, 16 calories of sugar 64 calories of honey

OpenStudy (sleepyjess):

this was a month ago.... -_-